# OmniSync: Towards Universal Lip Synchronization via Diffusion Transformers

### v1

@@ 1 Introduction
Traditional lip synchronization approaches rely heavily on reference frames combined with masked-frame inpainting prajwal2020lip ; zhang2023dinet ; guan2023stylesync ; guan2024resyncer . This methodology extracts appearance information from reference frames to inpaint masked regions in target frames—a process that introduces several critical limitations. These methods struggle with head pose variations, identity preservation, and artifact elimination, especially when target poses differ significantly from references peng2024synctalk .
@@ 3.2 Mask-Free Training Paradigm
where $x_{t}$ represents the noised version of target video $V_{ab}$ at timestep $t$, and DiT denotes our diffusion transformer, which predicts the denoised state at timestep $t-1$.
The standard diffusion loss used for training is defined as:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\epsilon,V_{cd},A_{ab},V_{ab}}[\|\epsilon-\epsilon_{\theta}(x_{t},V_{cd},A_{ab},t)\|_{2}^{2}],$ (2)
where $\epsilon$ denotes the random noise added during the forward diffusion process, and $\epsilon_{\theta}$ is the noise prediction network parameterized by $\theta$.
@@ 3.3 Progressive Noise Initialization
Standard diffusion-based generation tian2024emo typically begins from random noise (timestep $T$) and progressively denoises toward the final output (timestep 00). However, this approach often results in subtle but noticeable pose misalignments between generated content and original video frames, creating undesirable boundary artifacts and compromising identity preservation.
@@
This initialization strategy provides two significant advantages. First, it bypasses the early stages of diffusion (from $T$ down to $t_{\text{start}}$) where general facial structure is formed. This ensures that head pose and global structure are directly inherited from the source frame. Second, it reduces computational requirements by performing denoising only for the remaining steps, from $t_{\text{start}}$ down to 00.
@@ 4.1 Experimental Settings
Implementation Details. We implement our OmniSync model using the Diffusion Transformer architecture. The model is trained on a combined dataset for 80,000 steps using AdamW optimizer loshchilov2017decoupled with a learning rate of 1e-5. Training is completed in 80 hours using 64 NVIDIA A100 GPUs with a batch size of 64. Audio features are extracted via a pre-trained Whisper encoder, and text conditioning utilizes a T5 encoder. Training employs the timestep-dependent sampling threshold $t_{\text{threshold}}=850$. The experimental results indicate that excessive thresholds induce significant misalignment while insufficient values will leak the original lip shape. During inference we adopt our flow-matching-based progressive noise initialization starting at $\tau=0.92$, followed by 50 denoising steps.

### v2

@@ 1 Introduction
Traditional lip synchronization approaches rely heavily on reference frames combined with masked-frame inpainting prajwal2020lip ; zhang2023dinet ; guan2023stylesync ; guan2024resyncer . This methodology extracts appearance information from reference frames to inpaint masked regions in target frames—a process that introduces several critical limitations. These methods struggle with head pose variations, identity preservation, and artifact elimination, especially when target poses differ significantly from references peng2024synctalk ; peng2025synctalk++ .
@@ 3.2 Mask-Free Training Paradigm
where $x_{t}$ represents the noised version of target video $V_{ab}$ at timestep $t$, and DiT denotes our diffusion transformer, which predicts the denoised state at timestep t1t-1.
The CFM loss used for training is defined as:
$\mathcal{L}_{CFM}(\theta)=\mathbb{E}_{t,x_{t},V_{cd},A_{ab},V_{ab}}\left[\|v_{\theta}(x_{t},V_{cd},A_{ab},t)-u_{t}(x_{t}|V_{ab})\|_{2}^{2}\right],$ (2)
where $v_{\theta}(x_{t},V_{cd},A_{ab},t)$ is the learned velocity field predicted by DiT with conditioning on input video $V_{cd}$ and target audio $A_{ab}$, $u_{t}(x_{t}|V_{ab})$ is the conditional velocity field typically defined as $u_{t}(x_{t}|V_{ab})=(V_{ab}-x_{t})/(1-t)$ for the linear interpolation path $x_{t}=(1-t)x_{0}+tV_{ab}$.
@@ 3.3 Progressive Noise Initialization
Standard diffusion-based generation tian2024emo typically begins from random noise (timestep $T$) and progressively denoises toward the final output (timestep 0). However, this approach often results in subtle but noticeable pose misalignments between generated content and original video frames, creating undesirable boundary artifacts and compromising identity preservation.
@@
This initialization strategy provides two significant advantages. First, it bypasses the early stages of diffusion (from $T$ down to $t_{\text{start}}$) where general facial structure is formed. This ensures that head pose and global structure are directly inherited from the source frame. Second, it reduces computational requirements by performing denoising only for the remaining steps, from $t_{\text{start}}$ down to 0.
@@ 4.1 Experimental Settings
Implementation Details. We implement our OmniSync model using the Diffusion Transformer architecture. The model is trained on a combined dataset for 80,000 steps using AdamW optimizer loshchilov2017decoupled with a learning rate of 1e-5. Training is completed in 80 hours using 64 NVIDIA A100 GPUs with a batch size of 64. Audio features are extracted via a pre-trained Whisper encoder, and text conditioning utilizes a T5 encoder. Training employs the timestep‑dependent sampling threshold $t_{\text{threshold}}=850$. The experimental results indicate that excessive thresholds induce significant misalignment while insufficient values will leak the original lip shape. During inference we adopt our flow‑matching‑based progressive noise initialization starting at $\tau=0.92$, followed by 50 denoising steps.